# Geometry and cohomology of compactified Deligne–Lusztig varieties

### v1

@@ 3.2. Higher direct images of the structure sheaf for birational morphisms
Definition 3.6.
For any $v\in S^{*}$ such that $v=P_{m}(s,t)$ for a braid relation (3.1, 3.2, 3.3) in $W$, we set $R_{st}(v):=P_{m}(t,s)$.
Proposition 3.7.
Let $w_{1},\ldots,w_{r}\in W$ such that their corresponding Schubert varieties $\overline{{B}w_{i}{B}}/{B}$ are smooth for all $i=1,\ldots,r$. Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P_{m}(s,t)$ for a braid relation in $W$. Then for all $k\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{k}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
@@
$H^{k}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\omega_{X_{2}}\right),$
for $X_{i}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ and $i=1,2$.
@@
By assumption, we have $w_{j}=R_{st}(w_{j})$ in $W$, so
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R_{st}(w_{j})},\ldots,\underline{w_{r}}).$
For $X_{i}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 2.3 that $X_{i}$ is smooth and projective. Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by 2.12 $X_{i}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
There are natural projection maps ${\rm pr}_{0,\ell(w_{i})}:\overline{O}^{S^{*}}(w_{i})\to O(\underline{w_{i}})$ for all $i=1,\ldots r$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R_{st}(w_{j}))\to O(\underline{w_{j}})$. Similarly, the projection to the 00-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, . . ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
[…]
@@
[…]
The maps $f$ and $f^{\prime}$ are $G^{F}$-equivariant proper birational morphisms. It follows from the proof of Zariski’s main theorem ([Har, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp. $\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$ (resp. $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R_{st}(w_{j})\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, applying [CR, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$. Hence the Leray spectral sequence associated to $\phi$ degenerates. Since $\phi$ is $G^{F}$-equivariant, the isomorphism $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ is $G^{F}$-equivariant, thus functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{k}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{k}\left(X_{2},\mathcal{O}_{X_{2}}\right),$
for all $k\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nir, Corollary 2.10]. ∎
4. $\mathbb{P}^{1}$-fibrations and Cyclic shifting
We construct $\mathbb{P}^{1}$-fibrations between compactified Deligne–Lusztig varieties and relate their cohomology for the structure sheaf and canonical sheaf. Moreover, we use these $\mathbb{P}^{1}$-fibrations to indirectly show that the partial Frobenius on compactified Deligne–Lusztig varieties does not affect the cohomology for the structure sheaf and canonical sheaf.
@@ 4.1. $\mathbb{P}^{1}$-fibrations over Deligne–Lusztig varieties
Lemma 4.1.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. Then there is a $G^{F}$-equivariant morphism of $k$-schemes
$\pi_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is Zariski-locally a $\mathbb{P}^{1}$-bundle.
@@
For any $t\in S$, the projection map $O(\underline{t},\underline{t})\to O(\underline{t})$ given by $(B_{0},B_{1},B_{2})\mapsto(B_{0},B_{2})$ is a $\mathbb{P}^{1}$-bundle. Recall that it is a property of the Bruhat decomposition that if $(B_{0},B_{1})\in O(t)$ and $(B_{1},B_{2})\in O(t)$ then $(B_{0},B_{2})\in O(\underline{t})$, cf. [DL, §1.2]. By setting $t=F(s)$ and base change to $O(\underline{w_{1}},\ldots,\underline{w_{r}})$, we get again a $\mathbb{P}^{1}$-bundle
${\rm pr}_{F(s)}:O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})\longrightarrow O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
The partial Frobenius map induces a morphism of $k$-schemes
$F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)}).$
As $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})$ is a closed subscheme of $O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)})$, we may compose ${\rm pr}$ and $F_{1}$. It follows from the construction that
${\rm pr}_{F(s)}(X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)},\underline{F(s)}))\subseteq{\rm pr}_{0,r+1}^{-1}(\Gamma_{F}),$
where ${\rm pr}_{0,r+1}$ is the projection map
${\rm pr}_{0,r+1}:O(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow G/B\times G/B.$
Thus the image of the composition ${\rm pr}_{F(s)}\circ F_{1}$ is contained in $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$.
We define
$\pi_{1}:={\rm pr}_{F(s)}\circ F_{1}:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
which is $G^{F}$-equivariant by construction.
For any $(B_{0},B_{1})\in O(\underline{s})$, we have $(FB_{0},FB_{1})\in O(\underline{F(s)})$. Recall that via the projection to the first $(r+2)$-entries $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ identifies with a closed subscheme of $(G/B)^{r+1}$. Under this identification, we have an embedding
@@
Since $F(s)\in S$, for any $(B_{0},B_{1})\in O(\underline{s})$, if there exists a $B^{\prime}\in G/B$ with $(B^{\prime},FB_{1})\in O(\underline{F(s)})$, then we automatically have $(B^{\prime},FB_{0})\in O(\underline{F(s)})$. Thus $\iota$ is surjective with inverse given by $((B_{0}^{\prime},B_{1}^{\prime}),(B_{1}^{\prime},\ldots,B_{r+1}^{\prime}))\mapsto(B_{0}^{\prime},\ldots,B_{r+1}^{\prime})$. Note that here we have $(B_{r+1}^{\prime},FB_{0}^{\prime})\in O(\underline{F(s)})$. This yields that $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is isomorphic to $O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ and that $\pi_{1}$ is the composition of the isomorphism $\iota$ with the projection
@@
which is a Zariski locally trivial $\mathbb{P}^{1}$-bundle with a trivialising section. ∎
Remark 4.2.
The reason for the fibre product of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ with $O(\underline{s})$ yielding a Deligne–Lusztig variety is that the orbit of pairs $(B_{0},B_{1})$ determines the orbit of $(FB_{0},FB_{1})$. This allows us to verify that the tuples of the Borel subgroups live in the fibre product of the graph of Frobenius and a product of orbits. In general, taking a fibre product of a Deligne-Lusztig variety with $O(\underline{s})$ does not necessarily yield another Deligne–Lusztig variety.
Lemma 4.3.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$. There are $G^{F}$-equivariant étale $\mathbb{P}^{1}$-bundles
$\pi_{2}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}}),$
and
$F\circ\pi_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}).$
@@
Define $\pi_{2}$ by
$(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(B_{r+1},FB_{1},\ldots,FB_{r+1}).$
This is equivariant under the diagonal $G^{F}$-action. Since $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$. By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ [DL, §1.2].
Let $Z$ be given by the fibre product
@@
where $F=F_{1}^{r+1}$ is the Frobenius map. Points in $Z$ are denoted as tuples
@@
$(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ such that $FB_{0}^{\prime}=B_{r+1}$, $FB_{1}^{\prime}=FB_{1},\ldots,FB_{r}^{\prime}=FB_{r},F^{2}B_{0}^{\prime}=FB_{r+1}$. Here the only condition for $B_{0}$ is that $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$.
@@
[…]
@@
The diagram (4.2) yields an injective morphism
$\iota:Z\longrightarrow O(\underline{s})\times_{G/B}X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})}).$
On the other hand, given $(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ and arbitrary $B_{0}\in G/B$ with $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$, we have $(B_{0},B_{1}^{\prime})\in O(\underline{F^{-1}(s)})$ and $(FB_{0},FB_{0}^{\prime})\in O(\underline{s})$. This gives
@@
and that $\iota$ is surjective. Hence $\iota$ is an isomorphism and $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle. Let $\{U_{\alpha}\}_{\alpha}$ be a Zariksi covering of $X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ over which $d_{1}$ trivialises. Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is surjective and étale, $\{F(U_{\alpha})\}_{\alpha}$ is an étale covering of $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ over which $\pi_{2}$ trivialises as a $\mathbb{P}^{1}$-bundle.

### v2

@@ 3.2. Higher direct images of the structure sheaf for birational morphisms
Definition 3.2.1.
For any $v\in S^{*}$ such that $v=P(s,t)$ for a braid relation (1) in $W$, we set $R(s,t):=P(t,s)$.
Proposition 3.2.2.
Let $w_{1},\ldots,w_{r}\in W$ such that $\overline{{B}w_{j}{B}}$ is smooth for all $j=1,\ldots,r$. Suppose for $j\in\{1,\ldots,r\}$, the reduced expression of $w_{j}$ is of the form $w_{j}=P(s,t)$ for a braid relation in $W$. Then for all $i\geq 0$, there are $G^{F}$-equivariant isomorphisms
$H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right)$
@@
$H^{i}\left(X_{1},\omega_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\omega_{X_{2}}\right),$
for $X_{n}$ $(n\in\{1,2\})$ being one of the $k$-schemes ${\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, where $R(s,t)$ replaces $w_{j}$.
@@
By assumption, we have $w_{j}=P(s,t)=R(s,t)$ in $W$, so replacing $w_{j}$ by $R(s,t)$ does not affect the following compactified Deligne–Lusztig variety
$X(\underline{w_{1}},\ldots,\underline{w_{r}})=X(\underline{w_{1}},\ldots,\underline{R(s,t)},\ldots,\underline{w_{r}}).$
Replace $w_{j}$ by $R(s,t)$ in the expression $w_{1}\cdots w_{r}\in S^{*}$ and denote the resulting element by $w_{1}\cdots R(s,t)\cdots w_{r}$. For $X_{n}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$, ${\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$, it follows from the assumptions and Lemma 3.1.3 that $X_{n}$ is smooth and projective. Note that we have $I:={\text{supp}}_{F}(w_{1}\cdots w_{r})={\text{supp}}_{F}(w_{1}\cdots R(s,t)\cdots w_{r})={\text{supp}}_{F}(w_{1},\ldots,w_{r})$, and by Proposition 2.3.3 $X_{n}$ has $|G^{F}/P_{I}^{F}|$ irreducible components.
There are natural projection maps ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(w_{j})\to O(\underline{w_{j}})$ and ${\rm pr}_{0,\ell(w_{j})}:\overline{O}^{S^{*}}(R(s,t))\to O(\underline{w_{j}})$ for all $j=1,\ldots r$. Similarly, the projection to the 00-th, $\ell(w_{1})$-th, $\ell(w_{1})+\ell(w_{2})$-th, . . ., and $\ell(w_{1})+\cdots+\ell(w_{r})$-th coordinate yields the vertical maps in the cartesian diagrams
${{\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}),}$$\scriptstyle{f}$
@@
${{\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}}$${\overline{O}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$${X(\underline{w_{1}},\ldots,\underline{w_{r}})}$${O(\underline{w_{1}},\ldots,\underline{w_{r}}).}$$\scriptstyle{f^{\prime}}$
The maps $f$ and $f^{\prime}$ are $G^{F}$-equivariant proper birational morphisms.
It follows from the proof of Zariski’s main theorem ([Hartshorne, Corollary 11.4]) that $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$, for $\phi=f$ (resp. $\phi=f^{\prime}$), $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots w_{r})}$ (resp. $X_{1}={\overline{X}^{S^{*}}(w_{1}\cdots R(s,t)\cdots w_{r})}$) and $X_{2}=X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Moreover, applying [CR11, Corollary 3.2.10] yields $R\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$. Hence the Leray spectral sequence associated to $\phi$ degenerates.
Since $\phi$ is $G^{F}$-equivariant, the isomorphism $\phi_{*}\mathcal{O}_{X_{1}}\cong\mathcal{O}_{X_{2}}$ is $G^{F}$-equivariant, thus functoriality of the associated Leray spectral sequence yields $G^{F}$-equivariant isomorphisms on cohomology groups
$H^{i}\left(X_{1},\mathcal{O}_{X_{1}}\right)\overset{\sim}{\longrightarrow}H^{i}\left(X_{2},\mathcal{O}_{X_{2}}\right),$
for all $i\geq 0$. The $G^{F}$-equivariant isomorphisms for the cohomology of the canonical sheaf follow from an equivariant version of Serre duality [Nironi, Corollary 2.10]. ∎
4. $\mathbb{P}^{1}$-fibrations and $F$-conjugacy classes
We construct $\mathbb{P}^{1}$-fibrations between compactified Deligne–Lusztig varieties and relate their cohomology for the structure sheaf and the canonical sheaf. Moreover, we use these $\mathbb{P}^{1}$-fibrations to show that pairs of compactified Deligne–Lusztig varieties of the form ${\overline{X}^{S^{*}}(sv)}$ and ${\overline{X}^{S^{*}}(vF(s))}$ have isomorphic cohomology groups for the the structure sheaf and the canonical sheaf.
@@ 4.1. $\mathbb{P}^{1}$-fibrations over Deligne–Lusztig varieties
Lemma 4.1.1.
Let $w_{1},\ldots,w_{r},w_{1}^{\prime},\ldots,w_{m}^{\prime}\in W$ and $s\in S$. Then the projection map
$\phi:X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s},\underline{s},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{s},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})$
given by
$(B_{0},\ldots,B_{r},B_{r+1},B_{r+2},\ldots,B_{r+m+1},FB_{0})\mapsto(B_{0},\ldots,B_{r},B_{r+2},\ldots,B_{r+m+1},FB_{0})$ is a $G^{F}$-equivariant Zariski $\mathbb{P}^{1}$-bundle.
If $F(s)\in S$, then the $G^{F}$-equivariant morphism of $k$-schemes
$\pi:X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
given by $(B_{0},\ldots,B_{r+1},FB_{0})\mapsto(B_{1},\ldots,B_{r+1},FB_{1})$, is a Zariski $\mathbb{P}^{1}$-bundle.
@@
For any $t\in S$, the projection map $O(\underline{t},\underline{t})\to O(\underline{t})$ given by $(B_{0},B_{1},B_{2})\mapsto(B_{0},B_{2})$ is a $\mathbb{P}^{1}$-bundle. Recall that it is a property of the Bruhat decomposition that if $(B_{0},B_{1})\in O(t)$ and $(B_{1},B_{2})\in O(t)$ then $(B_{0},B_{2})\in O(\underline{t})$, cf. [DL, §1.2].
For $t=s$, we see that $\phi$ is the base change of $O(\underline{t},\underline{t})\to O(\underline{t})$ along the projection map $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{t},\underline{w_{1}^{\prime}},\ldots,\underline{w_{m}^{\prime}})\to O(\underline{t})$ given by $(B_{0}^{\prime},\ldots,FB_{0}^{\prime})\mapsto(B_{r}^{\prime},B_{r+1}^{\prime})$. The first claim follows.
Assume $F(s)\in S$ and consider $t=F(s)$. One can verify that $\pi$ is $G^{F}$-equivariant by its definition. Let $(B_{0},B_{1})\in O(\underline{s})$, so $(FB_{0},FB_{1})\in O(\underline{F(s)})$. Recall that via the projection to the first $(r+2)$-entries, $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ can be identified with a closed subscheme of $(G/B)^{r+2}$. Under this identification, we have an embedding
@@
Since $F(s)\in S$, for any $(B_{0},B_{1})\in O(\underline{s})$, if there exists a $B^{\prime}\in G/B$ with $(B^{\prime},FB_{1})\in O(\underline{F(s)})$, then we automatically have $(B^{\prime},FB_{0})\in O(\underline{F(s)})$. Thus $\iota$ is surjective with inverse given by $((B_{0}^{\prime},B_{1}^{\prime}),(B_{1}^{\prime},\ldots,B_{r+1}^{\prime}))\mapsto(B_{0}^{\prime},\ldots,B_{r+1}^{\prime})$. Therefore $X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ is isomorphic to $O(\underline{s})\times_{G/B}X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$. It follows that $\pi$ is the composition of the isomorphism $\iota$ with the projection
@@
which is a Zariski $\mathbb{P}^{1}$-bundle. ∎
Remark 4.1.2.
The reason for the fibre product of $X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)})$ with $O(\underline{s})$ to yield a Deligne–Lusztig variety is that the orbit of pairs $(B_{0},B_{1})$ determines the orbit of $(FB_{0},FB_{1})$. This allows us to verify that the tuples $(B_{0},\ldots,B_{r+1},FB_{0})$ live in the fibre product of the graph of $F$ and a product of orbits. In general, taking a fibre product of a Deligne-Lusztig variety with $O(\underline{s})$ does not necessarily yield another Deligne–Lusztig variety.
Lemma 4.1.3.
Let $w_{1},\ldots,w_{r}\in W$ and $s\in S$ such that $F(s)\in S$ and $\overline{Bw_{n}B}$ is smooth for $n=1,\ldots,r$. The morphisms
$\tau_{2}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}}),$
given by $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(B_{r+1},FB_{1},\ldots,FB_{r+1})$, and
$\tau_{1}:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})},\underline{s})\longrightarrow X(\underline{w_{1}},\ldots,\underline{w_{r}},\underline{F(s)}),$
given by $(B_{0},B_{1},\ldots,B_{r+1},FB_{0})\longmapsto(FB_{1},FB_{2}\ldots,F^{2}B_{1})$, are $G^{F}$-equivariant fppf $\mathbb{P}^{1}$-bundles.
@@
Given $(B_{0},B_{1})\in O(\underline{s})$, we know that $(FB_{0},FB_{1})\in O(F(\underline{s}))$. By construction we have $(B_{r+1},FB_{0})\in O(F(\underline{s}))$, so $(B_{r+1},FB_{1})\in O(F(\underline{s}))$ and $(FB_{r+1},F^{2}B_{1})\in O(F(\underline{s}))$ [DL, §1.2]. Thus $\tau_{1}$ and $\tau_{2}$ are well-defined. They are $G^{F}$-equivariant by construction.
Let $Z$ be the fibre product
@@
Geometric points in $Z$ are denoted as tuples
@@
$(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ such that $FB_{0}^{\prime}=B_{r+1}$, $FB_{1}^{\prime}=FB_{1},\ldots,FB_{r}^{\prime}=FB_{r},F^{2}B_{0}^{\prime}=FB_{r+1}$. Here the only condition on $B_{0}$ is that $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$.
@@
${Z}$${O(\underline{F^{-1}(s)})}$${X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})}$${G/B,}$$\scriptstyle{r_{2}}$$\scriptstyle{d_{1}}$$\scriptstyle{{\rm pr}_{1}}$$\scriptstyle{{\rm pr}_{0}}$ (4)
@@
The diagram (4) yields a morphism
$\iota:Z\longrightarrow O(\underline{F^{-1}(s)})\times_{G/B}X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})}).$
Given $(B_{0}^{\prime},B_{1}^{\prime},\ldots,B_{r}^{\prime},FB_{0}^{\prime})\in X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})$ and arbitrary $B_{0}\in G/B$ with $(B_{0},B_{0}^{\prime})\in O(\underline{F^{-1}(s)})$, we have $(B_{0},B_{1}^{\prime})\in O(\underline{F^{-1}(s)})$ and $(FB_{0},FB_{0}^{\prime})\in O(\underline{s})$. This gives
@@
and that $\iota$ is surjective. By the Bruhat decomposition, for every $i=1,\ldots,r$, there exists $v_{i}\in W$ such that $(B_{i},B_{i}^{\prime})\in O(v_{i})$. This implies $(FB_{i},FB_{i}^{\prime})\in O(F(v_{i}))$. We know $FB_{i}=FB_{i}^{\prime}$ by assumption, so $F(v_{i})=e$ is the neutral element in $W$ and $v_{i}=e$ because $F:W\to W$ is an automorphism, cf. [DL, §1.2]. It follows that $\iota$ is also injective and an isomorphism.
Hence $d_{1}$ is a Zariski $\mathbb{P}^{1}$-bundle. Since $F:X(\underline{F^{-1}(s)},\underline{F^{-1}(w_{1})},\ldots,\underline{F^{-1}(w_{r})})\to X(\underline{s},\underline{w_{1}},\ldots,\underline{w_{r}})$ is faithfully flat by our assumption on smoothness [GoerWedhAlgGeom1, Corollary 14.129], we deduce that $\tau_{2}$ is faithfully flat by faithfully flat descent [GoerWedhAlgGeom1, Corollary 14.12, Proposition 14.3(2)].